The volume of a cube is given by V = s3. Find the length of a side of a cube if the Volume is 729 cm3.
Answer by Cintchr(475)

find a # that multiplied by itself (three times) is equal to 729

5*5*5 = 125
6*6*6 = 216
7*7*7 = 343
8*8*8 = 512
9*9*9 = 729 BINGO!!

the cube is 9 x 9 x 9